Calculate the area expansivity and volume expansivity of a brass of area 120m and 120m respectively that assumes a new area and volume of 130m and 130m respectively when the linear expansivity is 0.00005 through a temperature of 50 degrees Celsius
1
Expert's answer
2020-06-18T10:59:31-0400
"\beta=2\alpha=2\frac{0.00005}{50}=2\cdot10^{-6}\degree C^{-1}"

"\gamma=3\alpha=3\frac{0.00005}{50}=3\cdot10^{-6}\degree C^{-1}"